# Next-to-leading-order QCD corrections to nucleon Dirac form factors

### v2

@@ Appendix A One-loop renormalization factor of non-local three-quark operators in KM scheme
$\left.+\Gamma_{000}\left\{\delta(x^{\prime}_{3}-x_{3})\left[\frac{4x_{2}}{x^{\prime}_{2}}\left[\frac{\theta(x_{1}-x^{\prime}_{1})}{x_{1}-x^{\prime}_{1}}\right]_{+}+\frac{4x_{1}}{x^{\prime}_{1}}\left[\frac{\theta(x_{2}-x^{\prime}_{2})}{x_{2}-x^{\prime}_{2}}\right]_{+}-\frac{4(x^{\prime}_{1}-x_{1})\theta(x_{2}-x^{\prime}_{2})}{x^{\prime}_{1}x^{\prime}_{2}}+\frac{4x_{2}}{x^{\prime}_{2}(1-x^{\prime}_{3})}\right]\right.\right.$
@@
The $+$-distribution is defined by
[…]
@@
$\frac{x_{2}}{x^{\prime}_{2}}\left[\frac{\theta(x_{1}-x^{\prime}_{1})}{x_{1}-x^{\prime}_{1}}\right]_{+}=\left[\frac{x_{2}}{x^{\prime}_{2}}\frac{\theta(x^{\prime}_{2}-x_{2})}{x^{\prime}_{2}-x_{2}}\right]_{+}-\delta(x_{1}-x^{\prime}_{1}),$ (27)
@@
${\cal Z}=\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}-\frac{\alpha_{s}}{12\pi}\frac{1}{\epsilon}\Bigg{\{}-6\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}$
@@
By projecting the renormalized three-quark operator onto $\Phi_{3}$ using (3), one readily reproduce the ERBL evolution kernel of $\Phi_{3}(x)$, whose explicit expression can be found in Eq. (4.18) of Passek-Kumericki:2008uqr .
Appendix B Analytical expressions for the convolution of the hard-scattering kernel
In this appendix, we enumerate the analytical expressions of various short-distance coefficients defined in (22), whose numerical values are tabulated in Table 1.
@@ Appendix B Analytical expressions for the convolution of the hard-scattering kernel
$+e_{d}\,d^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}(d^{ijkl}_{1}L_{\mu}+d^{ijkl}_{2})\right]+{\cal O}(\alpha_{s}^{2})\Big{\}}.$
@@
    $c^{0000}_{0}=1,\quad d^{0000}_{0}=2,\quad c^{0000}_{1}=d_{1}=\frac{29}{6}$
    $c_{2}^{0000}=-92\zeta_{5}+\frac{3719}{30}\zeta_{3}-\frac{1897}{60},$
    $d_{2}^{0000}=\frac{3}{2}\zeta_{5}+\frac{209}{120}\zeta_{3}+\frac{1741}{120}.$ (31)
@@
    $c^{1000}_{0}=\frac{49}{6},d^{1000}_{0}=-\frac{49}{6},\quad c^{1000}_{1}=d_{1}=\frac{97}{18}$
    $c_{2}^{1000}=d_{2}^{1000}=-100\zeta_{5}+\frac{3203}{40}\zeta_{3}+\frac{981829}{30240}.$ (32)
@@
    $c_{0}^{1010}=\frac{1519}{9},\quad d_{0}^{1010}=\frac{1568}{9},\quad c^{1010}_{1}=d^{1010}_{1}=\frac{107}{18}$
    $c_{2}^{1010}=-\frac{55926\zeta_{5}}{217}+\frac{151409\zeta_{3}}{868}+\frac{11652817}{133920},$
    $d_{2}^{1010}=-\frac{22419\zeta_{5}}{224}+\frac{160177\zeta_{3}}{2240}+\frac{5656271}{120960}.$ (33)
@@
    $c_{0}^{1100}=\frac{7}{2},\quad d_{0}^{1100}=-\frac{7}{2},\quad c^{1100}_{1}=d^{1100}_{1}=\frac{11}{2}$
    $c_{2}^{1100}=-\frac{130}{3}\zeta_{5}+\frac{5161}{72}\zeta_{3}-\frac{369}{32},$
    $d_{2}^{1100}=\frac{92}{3}\zeta_{5}-\frac{11953}{360}\zeta_{3}+\frac{38417}{1440}.$ (34)
@@
    $c_{0}^{1110}=\frac{343}{9},\quad d_{0}^{1110}=-\frac{343}{9},\quad c^{1110}_{1}=d^{1110}_{1}=\frac{109}{18}$
    $c_{2}^{1110}=d_{2}^{1110}=-\frac{2400\zeta_{5}}{7}+\frac{447161\zeta_{3}}{1960}+\frac{23648419}{211680}.$
@@
    $c_{0}^{1111}=\frac{539}{9},\quad d_{0}^{1111}=\frac{196}{9},\quad c^{1111}_{1}=d^{1111}_{1}=\frac{37}{6}$
    $c_{2}^{1111}=\frac{762}{11}\zeta_{5}-\frac{19183}{770}\zeta_{3}-\frac{4745929}{332640},$
    $d_{2}^{1111}=-\frac{141}{4}\zeta_{5}+\frac{17807}{560}\zeta_{3}+\frac{750391}{30240}.$ (36)
@@
    $c_{0}^{2000}=\frac{427}{20},\quad d_{0}^{2000}=\frac{161}{20},\quad c^{2000}_{1}=d^{2000}_{1}=\frac{103}{18}$
    $c_{2}^{2000}=-\frac{74566}{427}\zeta_{5}+\frac{18937349}{153720}\zeta_{3}+\frac{110572513}{1844640},$
    $d_{2}^{2000}=\frac{1780}{161}\zeta_{5}-\frac{244199}{57960}\zeta_{3}+\frac{12993881}{695520}.$ (37)
@@
    $c_{0}^{2010}=\frac{539}{6},\quad d_{0}^{2010}=-\frac{539}{6},\quad c^{2010}_{1}=d^{2010}_{1}=\frac{113}{18}$
    $c_{2}^{2010}=d_{2}^{2010}=-\frac{119124}{385}\zeta_{5}+\frac{1590773}{7700}\zeta_{3}+\frac{4637387}{44352}.$
@@
    $c_{0}^{2011}=\frac{931}{10},\quad d_{0}^{2011}=\frac{49}{5},\quad c^{2011}_{1}=d^{2011}_{1}=\frac{115}{18}$
    $c_{2}^{2011}=-\frac{11792}{19}\zeta_{5}+\frac{19751197}{47880}\zeta_{3}+\frac{622684033}{3447360},$
    $d_{2}^{2011}=-\frac{7471}{14}\zeta_{5}+\frac{367657}{1008}\zeta_{3}+\frac{55780799}{362880}.$ (39)
@@
    $c_{0}^{2020}=\frac{23471}{100},\quad d_{0}^{2020}=\frac{1372}{25},\quad c^{2020}_{1}=d^{2020}_{1}=\frac{119}{18}$
    $c_{2}^{2020}=-\frac{3537990}{3353}\zeta_{5}+\frac{46755193}{67060}\zeta_{3}+\frac{702816409}{2414160},$
    $d_{2}^{2020}=\frac{430209}{784}\zeta_{5}-\frac{2824781}{7840}\zeta_{3}-\frac{16409957}{161280}.$ (40)
@@
    $c_{0}^{2100}=\frac{35}{4},\quad d_{0}^{2100}=-\frac{35}{4},\quad c^{2100}_{1}=d^{2100}_{1}=\frac{107}{18}$
    $c_{2}^{2100}=d_{2}^{2100}=\frac{2732}{35}\zeta_{5}-\frac{179843}{4200}\zeta_{3}-\frac{7131}{5600}.$ (41)
@@
    $c_{0}^{2110}=d_{0}^{2110}=\frac{637}{4},\quad c^{2110}_{1}=d^{2110}_{1}=\frac{13}{2}$
    $c_{2}^{2110}=\frac{4908}{13}\zeta_{5}-\frac{447763}{1820}\zeta_{3}-\frac{49062911}{786240},$
    $d_{2}^{2110}=-\frac{9546}{91}\zeta_{5}+\frac{129473}{1820}\zeta_{3}+\frac{6383213}{112320}.$ (42)
@@
    $c_{0}^{2111}=\frac{539}{12},\quad d_{0}^{2111}=-\frac{539}{12},\quad c^{2111}_{1}=d^{2111}_{1}=\frac{119}{18}$
    $c_{2}^{2111}=d_{2}^{2111}=\frac{47940}{77}\zeta_{5}-\frac{307441}{770}\zeta_{3}-\frac{88027997}{665280}.$
@@
    $c_{0}^{2120}=\frac{3381}{40},\quad d_{0}^{2120}=-\frac{3381}{40},\quad c^{2120}_{1}=d^{2120}_{1}=\frac{41}{6}$
    $c_{2}^{2120}=d_{2}^{2120}=\frac{36348}{161}\zeta_{5}-\frac{1422557}{9660}\zeta_{3}-\frac{85715459}{4173120}.$
@@
    $c_{0}^{2121}=d_{0}^{2121}=\frac{637}{4},\quad c^{2121}_{1}=d^{2121}_{1}=\frac{127}{18}$
    $c_{2}^{2121}=-\frac{94014}{91}\zeta_{5}+\frac{124647}{182}\zeta_{3}+\frac{675228271}{2358720},$
    $d_{2}^{2121}=\frac{57285}{91}\zeta_{5}-\frac{106337}{260}\zeta_{3}-\frac{32207713}{262080}.$ (45)
@@
    $c_{0}^{2200}=\frac{16}{5},\quad d_{0}^{2200}=\frac{1}{10},\quad c^{2200}_{1}=d^{2200}_{1}=6$
    $c_{2}^{2200}=-\frac{3543}{16}\zeta_{5}+\frac{1924849}{11520}\zeta_{3}+\frac{7842193}{138240},$
    $d_{2}^{2200}=840\zeta_{5}-\frac{144079}{360}\zeta_{3}-\frac{1479289}{4320}.$ (46)
@@
    $c_{0}^{2210}=\frac{49}{4},\quad d_{0}^{2210}=-\frac{49}{4},\quad c^{2210}_{1}=d^{2210}_{1}=\frac{59}{9}$
    $c_{2}^{2210}=d_{2}^{2210}=\frac{34068}{35}\zeta_{5}-\frac{109994}{175}\zeta_{3}-\frac{4467161}{20160}.$
@@
    $c_{0}^{2211}=\frac{427}{60},\quad d_{0}^{2211}=\frac{161}{60},\quad c^{2211}_{1}=d^{2211}_{1}=\frac{20}{3}$
    $c_{2}^{2211}=-\frac{180108}{61}\zeta_{5}+\frac{1752946}{915}\zeta_{3}+\frac{140407409}{175680},$
    $d_{2}^{2211}=-\frac{8352}{23}\zeta_{5}+\frac{15283}{69}\zeta_{3}+\frac{9688567}{66240}.$ (48)
@@
    $c_{0}^{2220}=\frac{6069}{200},\quad d_{0}^{2220}=\frac{987}{200},\quad c^{2220}_{1}=d^{2220}_{1}=\frac{62}{9}$
    $c_{2}^{2220}=\frac{4267860}{2023}\zeta_{5}-\frac{28050761}{20230}\zeta_{3}-\frac{25391852771}{52436160},$
    $d_{2}^{2220}=\frac{1066224}{329}\zeta_{5}-\frac{7005001}{3290}\zeta_{3}-\frac{6511339837}{8527680}.$ (49)
@@
    $c_{0}^{2221}=\frac{287}{40},\quad d_{0}^{2221}=-\frac{287}{40},\quad c^{2221}_{1}=d^{2221}_{1}=\frac{64}{9}$
    $c_{2}^{2221}=d_{2}^{2221}=-\frac{2540124}{287}\zeta_{5}+\frac{16718021}{2870}\zeta_{3}+\frac{5492284859}{2479680}.$
@@
    $c_{0}^{2222}=\frac{779}{100},\quad d_{0}^{2222}=\frac{43}{25},\quad c^{2222}_{1}=d^{2222}_{1}=\frac{43}{6}$
    $c_{2}^{2222}=\frac{3727080}{779}\zeta_{5}-\frac{12158634}{3895}\zeta_{3}-\frac{11846765099}{10095840},$
    $d_{2}^{2222}=-\frac{20277}{43}\zeta_{5}+\frac{571137}{1720}\zeta_{3}+\frac{278914037}{2229120}.$ (51)

### v3

@@ A.1 One-loop renormalization factor of non-local three-quark operators in KM scheme
[…]
@@
The ++-distribution is defined by
$\int_{0}^{1}dx_{i}\left[\frac{\theta(x_{i}-x^{\prime}_{i})}{x_{i}-x^{\prime}_{i}}\right]_{+}f(x_{i})=\int_{0}^{1}dx_{i}\frac{\theta(x_{i}-x^{\prime}_{i})}{x_{i}-x^{\prime}_{i}}\left[f(x_{i})-f(x^{\prime}_{i})\right],$ (18)
@@
$\frac{x_{2}}{x^{\prime}_{2}}\left[\frac{\theta(x_{1}-x^{\prime}_{1})}{x_{1}-x^{\prime}_{1}}\right]_{+}=\left[\frac{x_{2}}{x^{\prime}_{2}}\frac{\theta(x^{\prime}_{2}-x_{2})}{x^{\prime}_{2}-x_{2}}\right]_{+}-\delta(x_{1}-x^{\prime}_{1}),$
@@
$\hskip-14.22636pt{\cal Z}=\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}-\frac{\alpha_{s}}{12\pi}\frac{1}{\epsilon}\Bigg{\{}-6\delta(x_{1}-x^{\prime}_{1})\delta(x_{2}-x^{\prime}_{2})\Gamma_{000}$
@@
By projecting the renormalized three-quark operator onto $\Phi_{3}$ using Eq.(3) of main text, one readily reproduce the ERBL evolution kernel of $\Phi_{3}(x)$, whose explicit expression can be found in Eq. (4.18) of Passek-Kumericki:2008uqr .
A.2 Analytical expressions of low-lying ${\cal H}^{ijkl}$
In this section, we enumerate the analytical expressions of various short-distance coefficients appearing in the nucleon form factor predictions:
@@ A.2 Analytical expressions of low-lying ${\cal H}^{ijkl}$
${\cal H}^{ijkl}$ $=$ $\frac{1600\pi^{2}\alpha_{s}^{2}}{3}\Big{\{}e_{u}\,c^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}\left(c^{ijkl}_{1}L_{\mu}+c^{ijkl}_{2}\right)\right]+e_{d}\,d^{ijkl}_{0}\left[1+{\alpha_{s}\over\pi}(d^{ijkl}_{1}L_{\mu}+d^{ijkl}_{2})\right]+{\cal O}(\alpha_{s}^{2})\Big{\}}.$ (22)
@@
    $c^{0000}_{0}=1,\;d^{0000}_{0}=2,\;c^{0000}_{1}=d^{0000}_{1}=\frac{29}{6},$
    $c_{2}^{0000}=-92\zeta_{5}+\frac{3719}{30}\zeta_{3}-\frac{1937}{60},\;d_{2}^{0000}=\frac{3}{2}\zeta_{5}+\frac{209}{120}\zeta_{3}+\frac{1661}{120}.$ (23)
@@
    $c^{1000}_{0}=\frac{49}{6},\;d^{1000}_{0}=-\frac{49}{6},\;c^{1000}_{1}=d^{1000}_{1}=\frac{97}{18},\;c_{2}^{1000}=d_{2}^{1000}=-100\zeta_{5}+\frac{3203}{40}\zeta_{3}+\frac{961189}{30240}.$ (24)
@@
    $c_{0}^{1010}=\frac{1519}{9},\;d_{0}^{1010}=\frac{1568}{9},\;c^{1010}_{1}=d^{1010}_{1}=\frac{107}{18},$
    $c_{2}^{1010}=-\frac{55926\zeta_{5}}{217}+\frac{151409\zeta_{3}}{868}+\frac{11541697}{133920},\;d_{2}^{1010}=-\frac{22419\zeta_{5}}{224}+\frac{160177\zeta_{3}}{2240}+\frac{5573321}{120960}.$ (25)
@@
    $c_{0}^{1100}=\frac{7}{2},\;d_{0}^{1100}=-\frac{7}{2},\;c^{1100}_{1}=d^{1100}_{1}=\frac{11}{2},$
    $c_{2}^{1100}=-\frac{130}{3}\zeta_{5}+\frac{5161}{72}\zeta_{3}-\frac{401}{32},\;d_{2}^{1100}=\frac{92}{3}\zeta_{5}-\frac{11953}{360}\zeta_{3}+\frac{37937}{1440}.$ (26)
@@
    $c_{0}^{1110}=\frac{343}{9},\;d_{0}^{1110}=-\frac{343}{9},\;c^{1110}_{1}=d^{1110}_{1}=\frac{109}{18},\;c_{2}^{1110}=d_{2}^{1110}=-\frac{2400\zeta_{5}}{7}+\frac{447161\zeta_{3}}{1960}+\frac{23482099}{211680}.$ (27)
@@
    $c_{0}^{1111}=\frac{539}{9},\;d_{0}^{1111}=\frac{196}{9},\;c^{1111}_{1}=d^{1111}_{1}=\frac{37}{6},$
    $c_{2}^{1111}=\frac{762}{11}\zeta_{5}-\frac{19183}{770}\zeta_{3}-\frac{4945849}{332640},\;d_{2}^{1111}=-\frac{141}{4}\zeta_{5}+\frac{17807}{560}\zeta_{3}+\frac{737371}{30240}.$ (28)
@@
    $c_{0}^{2000}=\frac{427}{20},\;d_{0}^{2000}=\frac{161}{20},\;c^{2000}_{1}=d^{2000}_{1}=\frac{103}{18},$
    $c_{2}^{2000}=-\frac{74566}{427}\zeta_{5}+\frac{18937349}{153720}\zeta_{3}+\frac{1791733}{30240},\;d_{2}^{2000}=\frac{1780}{161}\zeta_{5}-\frac{244199}{57960}\zeta_{3}+\frac{12557081}{695520}.$ (29)
@@
    $c_{0}^{2010}=\frac{539}{6},\;d_{0}^{2010}=-\frac{539}{6},\;c^{2010}_{1}=d^{2010}_{1}=\frac{113}{18},\;c_{2}^{2010}=d_{2}^{2010}=-\frac{119124}{385}\zeta_{5}+\frac{1590773}{7700}\zeta_{3}+\frac{23033831}{221760}.$ (30)
@@
    $c_{0}^{2011}=\frac{931}{10},\;d_{0}^{2011}=\frac{49}{5},\;c^{2011}_{1}=d^{2011}_{1}=\frac{115}{18},$
    $c_{2}^{2011}=-\frac{11792}{19}\zeta_{5}+\frac{19751197}{47880}\zeta_{3}+\frac{619980913}{3447360},\;d_{2}^{2011}=-\frac{7471}{14}\zeta_{5}+\frac{367657}{1008}\zeta_{3}+\frac{55349039}{362880}.$ (31)
@@
    $c_{0}^{2020}=\frac{23471}{100},\;d_{0}^{2020}=\frac{1372}{25},\;c^{2020}_{1}=d^{2020}_{1}=\frac{119}{18},$
    $c_{2}^{2020}=-\frac{3537990}{3353}\zeta_{5}+\frac{46755193}{67060}\zeta_{3}+\frac{701158249}{2414160},\;d_{2}^{2020}=\frac{430209}{784}\zeta_{5}-\frac{2824781}{7840}\zeta_{3}-\frac{16515077}{161280}.$ (32)
@@
    $c_{0}^{2100}=\frac{35}{4},\;d_{0}^{2100}=-\frac{35}{4},\;c^{2100}_{1}=d^{2100}_{1}=\frac{107}{18},\;c_{2}^{2100}=d_{2}^{2100}=\frac{2732}{35}\zeta_{5}-\frac{179843}{4200}\zeta_{3}-\frac{92179}{50400}.$ (33)
@@
    $c_{0}^{2110}=d_{0}^{2110}=\frac{637}{4},\;c^{2110}_{1}=d^{2110}_{1}=\frac{13}{2},$
    $c_{2}^{2110}=\frac{4908}{13}\zeta_{5}-\frac{447763}{1820}\zeta_{3}-\frac{49524911}{786240},\;d_{2}^{2110}=-\frac{9546}{91}\zeta_{5}+\frac{129473}{1820}\zeta_{3}+\frac{6308573}{112320}.$ (34)
@@
    $\;c_{0}^{2111}=\frac{539}{12},\;d_{0}^{2111}=-\frac{539}{12},\;c^{2111}_{1}=d^{2111}_{1}=\frac{119}{18},\;c_{2}^{2111}=d_{2}^{2111}=\frac{47940}{77}\zeta_{5}-\frac{307441}{770}\zeta_{3}-\frac{88449677}{665280}.$ (35)
@@
    $c_{0}^{2120}=\frac{3381}{40},\;d_{0}^{2120}=-\frac{3381}{40},\;c^{2120}_{1}=d^{2120}_{1}=\frac{41}{6},\;c_{2}^{2120}=d_{2}^{2120}=\frac{36348}{161}\zeta_{5}-\frac{1422557}{9660}\zeta_{3}-\frac{88420259}{4173120}.$ (36)
@@
    $c_{0}^{2121}=d_{0}^{2121}=\frac{637}{4},\;c^{2121}_{1}=d^{2121}_{1}=\frac{127}{18},$
    $c_{2}^{2121}=-\frac{94014}{91}\zeta_{5}+\frac{124647}{182}\zeta_{3}+\frac{673746511}{2358720},\;d_{2}^{2121}=\frac{57285}{91}\zeta_{5}-\frac{106337}{260}\zeta_{3}-\frac{2490181}{20160}.$ (37)
@@
    $c_{0}^{2200}=\frac{16}{5},\;d_{0}^{2200}=\frac{1}{10},\;c^{2200}_{1}=d^{2200}_{1}=6,$
    $c_{2}^{2200}=-\frac{3543}{16}\zeta_{5}+\frac{1924849}{11520}\zeta_{3}+\frac{7761793}{138240},\;d_{2}^{2200}=840\zeta_{5}-\frac{144079}{360}\zeta_{3}-\frac{1488889}{4320}.$ (38)
@@
    $c_{0}^{2210}=\frac{49}{4},\;d_{0}^{2210}=-\frac{49}{4},\;c^{2210}_{1}=d^{2210}_{1}=\frac{59}{9},\;c_{2}^{2210}=d_{2}^{2210}=\frac{34068}{35}\zeta_{5}-\frac{109994}{175}\zeta_{3}-\frac{4478633}{20160}.$ (39)
@@
    $c_{0}^{2211}=\frac{427}{60},\;d_{0}^{2211}=\frac{161}{60},\;c^{2211}_{1}=d^{2211}_{1}=\frac{20}{3},$
    $c_{2}^{2211}=-\frac{180108}{61}\zeta_{5}+\frac{1752946}{915}\zeta_{3}+\frac{46735883}{58560},\;d_{2}^{2211}=-\frac{8352}{23}\zeta_{5}+\frac{15283}{69}\zeta_{3}+\frac{3215629}{22080}.$ (40)
@@
    $c_{0}^{2220}=\frac{6069}{200},\;d_{0}^{2220}=\frac{987}{200},\;c^{2220}_{1}=d^{2220}_{1}=\frac{62}{9},$
    $c_{2}^{2220}=\frac{4267860}{2023}\zeta_{5}-\frac{28050761}{20230}\zeta_{3}-\frac{25422835331}{52436160},\;d_{2}^{2220}=\frac{1066224}{329}\zeta_{5}-\frac{7005001}{3290}\zeta_{3}-\frac{6517209757}{8527680}.$ (41)
@@
    $\hskip-7.11317ptc_{0}^{2221}=\frac{287}{40},\;d_{0}^{2221}=-\frac{287}{40},\;c^{2221}_{1}=d^{2221}_{1}=\frac{64}{9},\;c_{2}^{2221}=d_{2}^{2221}=-\frac{2540124}{287}\zeta_{5}+\frac{16718021}{2870}\zeta_{3}+\frac{5490319259}{2479680}.$ (42)
@@
    $c_{0}^{2222}=\frac{779}{100},\;d_{0}^{2222}=\frac{43}{25},\;c^{2222}_{1}=d^{2222}_{1}=\frac{43}{6},$
    $c_{2}^{2222}=\frac{3727080}{779}\zeta_{5}-\frac{12158634}{3895}\zeta_{3}-\frac{11850532139}{10095840},\;d_{2}^{2222}=-\frac{20277}{43}\zeta_{5}+\frac{571137}{1720}\zeta_{3}+\frac{278183957}{2229120}.$ (43)